OpenStudy (idealist10):

Find dy/dx and d^2*y/dx^2 if x=integral of (sin u)/u du from 1 to t and y=integral of e^u du from 2 to ln t.

OpenStudy (idealist10):

\[x=\int\limits_{1}^{t}\frac{ \sin u }{ u }du\]

OpenStudy (idealist10):

\[y=\int\limits_{2}^{\ln t}e^u du\]

OpenStudy (idealist10):

@zepdrix

OpenStudy (aum):

\[ \frac{dy}{dx} = \frac{dy}{dt} * \frac{dt}{dx} = \frac{dy}{dt} \div \frac{dx}{dt} \]

OpenStudy (aum):

\[ y=\int\limits_{2}^{\ln t}e^u du \\ \frac{dy}{dt} = e^{\ln(t)} * \frac{d}{dt} \ln(t) = \frac{e^{\ln(t)} } { t } = \frac{t} { t } = 1\\ \text{ } \\ x=\int\limits_{1}^{t}\frac{ \sin (u) }{ u }du \\ \frac{dx}{dt} = \frac{ \sin(t) }{ t } \\ \frac{dy}{dx} = \frac{dy}{dt} \div \frac{dx}{dt} ~~~ \text{(provided } \frac{dx}{dt} \ne 0) \\ \frac{dy}{dx}= 1 \div \frac{\sin(t) }{t} \\ \frac{dy}{dx}= \frac{t }{\sin(t)} \\ \]

OpenStudy (aum):

\[ \frac{d^2y}{dx^2} = \frac{d}{dx}\left(\frac{dy}{dx}\right) = \frac{d}{dt}\left(\frac{dy}{dx}\right) * \frac{dt}{dx} = \frac{d}{dt}\left(\frac{dy}{dx}\right) \div \frac{dx}{dt} = ? \]

OpenStudy (idealist10):

Thanks a lot!